# Conditional [MASK] Discrete Diffusion Language Model

### v3

@@ 3.2 DDLM with Entropy-based Denoising
where $p_{\theta}(x^{\prime}_{i};X^{(t)})$ is the softmax probability of token $x^{\prime}_{i}$ at position $i$ in sequence $X^{(t)}$, and $t$ denotes the diffusion timestep.
@@ 5.1 Tasks & Details
We conduct experiments on various conditional generation datasets. Detailed explanation of the conditional generation datasets are in Appendix F.1. In particular, we focus on two datasets that significantly differ in their level of conditional constraints: RocStories (Mostafazadeh et al., 2016), which is relatively open-ended, and Paradetox (Logacheva et al., 2022), which imposes the strongest conditional constraints. We select the conditional dataset that GPT-2 faces in generating sentences of appropriate length under specified conditional constraints. The maximum lengths of Paradetox and RocStories is set to 64, based on data statistics, and other details are in Appendix F. The number of steps of our model is configured to 5 with a naive categorical sampling with a sample size of 20 and select final 5 samples based on Perplexity score. We use 1 A100 GPU with the batch size as 256.
@@ 5.2 Baselines
For ARMs (Vaswani et al., 2023), we employ GPT-2 (Radford et al., 2019) and GPT-3.5-turbo with four-shot prompt. More experimental details of GPT-3.5 can be found in Appendix J, and H. For CMLMs, we utilize CMLM-Mask-Predict (Ghazvininejad et al., 2019a) and DisCo-Easy-First (Kasai et al., 2020), which are transformer-based NAR models. For CDLMs, our baseline includes DiffuSeq (Gong et al., 2023a), LD4LG (Lovelace et al., 2023), and DINOISER (Ye et al., 2024). DiffuSeq and DINOISER is designed for sequence-to-sequence applications, and LD4LG adopt BART as denoising init point. For DDLMs, we utilize DiffusionBERT (He et al., 2022), applying BERT into DDLMs, and SEDD (Lou et al., 2024), a powerful open-ended generation DDLM. For SEDD, we download the pre-trained version and fine-tune it.
@@ 5.3 Metrics
In addition to our theoretically guided methods, we evaluate performance using multiple metrics. Specifically, we use Perplexity (PPL) based-on GPT-2 Large and GPT-2 XL as an automated metric, MAUVE Pillutla et al. (2021) to assess style consistency between the training data and generated outputs, SOME (Yoshimura et al., 2020) to score the grammar, Mean Opinion Score (MOS) from human evaluations to gauge text quality, and LLM score such as LLM-c (Lin and Chen, 2023) to measure the plausibility of the narratives as a sub-metric. Further details on LLM scores and other evaluation settings are provided in Appendix F.3 and H.1.
@@
Following our quality assessment, we evaluate diversity through three different measures: an automatic frequency-based metric n-gram Vendi Score(VS n-gram) Friedman and Dieng (2023), a neural network–based semantic metric SimCSE Vendi Score (VS emb), and a human evaluation score MOS. More details related to MOS and other scores are in Appendix F.3.
6 Results
@@
Overall, Diffusion-EAGS consistently demonstrates strong performance across diverse conditional generation tasks, combining low perplexity and high human evaluation scores with the ability to generate richly varied text.
@@ F.2 Dataset Explanations
Deontology The objective of Deontology (Hendrycks et al., 2023) is to to evaluate the capability of models to make ethical judgments from a deontological perspective. The dataset contains scenarios focusing on interpersonal dynamics and everyday occurrences.
@@ F.3 Experimental Details
We employ roberta-base as MLM with learning rate 5e-4. The maximum lengths for QG, QQP, and Paradetox is set to 64, while for Deontology and DialogSum set to 48 and 292, respectively, based on data statistics. We test 20 conditions with 5 outputs in total 100, which is not used for training. The number of steps is configured to 5. We then perform a naive categorical sampling with a sample size of 20 and select final 5 samples based on PPL. We use 1 A100 GPU with the batch size as 256. Experimental details of LLMs are in Appendix J, machine translation in Appendix H.
@@
To measure the quality of the generated texts, we use Perplexity based-on GPT-2 Large and GPT-2 XL, SOME (Yoshimura et al., 2020), the grammar metric based on corpus, LLM-c (Lin and Chen, 2023) to measure the plausibility of the narratives, LLM-t (Koh et al., 2024a) to measure toxicity, and MAUVE Pillutla et al. (2021), measuring a reflectiveness of training dataset characteristics of generate outputs. MAUVE score of 1 indicates that the output perfectly matches the training dataset as a neural database. For Mean Opinion Score (MOS), we get 5 outputs from each condition. Subsequently, hired four integrated ph.d course work annotators in the university NLP research lab evaluate the generated text based on two criteria: (1) semantic reflectiveness of the condition, indicating how accurately the condition is represented in the text, and (2) sentence completeness, assessing overall grammatical and semantic coherence. Each criterion was rated on a scale from 0 to 1. Subsequently, these scores are normalized and averaged to obtain a final score ranging from 0 to 1. In our evaluation, Fleiss’ kappa Fleiss (1971) is exceeded 0.7 as assessing sentence quality is both intuitive and relatively non-controversial among the annotators.
@@ G.1 Fine-Grained Comparison
Diffusion-EAGS demonstrates the highest MAUVE score of 0.733 in Table 8-ParaDetox, and high level of text quality surpassing that of GPT-2 in Table 9 in text quality. ParaDetox is colloquial dataset including slang, numerous abbreviations, and various perturbations, so our model demonstrate robustness to such perturbations with 69.5 PPL. As for diversity, our model consistently outperforms GPT models in VS(ngram) and VS(emb) in Table 28, and 9.
Notably, CDLMs demonstrate a noticeable deficiency in diversity. In contrast, our model excels at producing significantly more diverse sequences. Furthermore, our models require only a few steps, while resulting in higher quality and diversity.
@@ G.2 Quality Recheck – LLM score & Dialogue Summarization
Ours 0.409 0.174 0.536 4.114 2.591 0.252 0.253 0.632
SEDD 0.179 0.032 0.999 4.216 2.576 0.211 0.200 0.609
DINOISER 0.209 0.031 0.337 1.247 1.227 0.926 0.256 0.633
Table 10: DialogueSum Experiment
@@
For Question Generation (QG), we employ the widely adopted GPT-as-a-Judge framework Zheng et al. (2023) to evaluate the quality of generations produced by our model and the baselines on the QG dataset. We adopt a pairwise evaluation setting, following the system and input prompts specified in Zheng et al. (2023) for the pairwise comparison. The factors specified to be evaluated are 1) coherency 2) grammatical correctness 3) semantic soundness 4) diversity and 5) being a more reasonable question to the input (condition) text. We employ the gpt-4 model. The result is as follows:
@@
Below, we report the Mean Opinion Score (MOS) averages and standard deviations (std) in the following order: DiffusionBERT, LD4LG, GPT-2, Dinoiser, and our method. First, the average scores of semantic reflection are 0.98, 0.90, 0.94, 0.98, and 0.97, respectively, with standard deviations of 0.14, 0.30, 0.24, 0.14, and 0.16. Second, the average scores of sentence completeness are 0.78, 0.92, 0.72, 0.84, and 0.90, respectively, with standard deviations of 0.18, 0.14, 0.28, 0.15, and 0.15. Third, average scores of diversity are 2, 1, 2.65, 1, and 4.6, respectively, with standard deviations of 1.3, 0, 1.45, 0, and 0.7. GPT-3.5-turbo’s std is almost 0.
@@ G.3 Machine Translation : Bilinguality & Low Resource Settings
Labeled datasets used in conditional generation tasks are typically limited in size and sometimes multilingual. To further assess our model’s performance in conditional generation, particularly in terms of language extension and resource scarcity, we conduct additional experiments on a translation task. We conduct additional experiments on CMLMs such as Mask-and-Predict and Easy-First, diffusion models such as Diffuseq-v2 (Gong et al., 2023b) and SEDD, traditional translation models such as mBART-50 (Tang et al., 2020) and NLLB. For evaluation metrics, we utilize sacreBLEU (Post, 2018) and neural-net scores such as COMET (Rei et al., 2020) and XCOMET (Guerreiro et al., 2023). More details are provided in Appendix L.
@@ G.4 Diversity Analysis
We summarize the generation trends of the models presented in Table above. We observe that when a fine-tuned GPT-2 is tasked with strongly constrained conditional generation, it struggles to properly terminate sentences with an <eos> token. In particular, it shows limitations when handling semantic leaps or clearly delineated structural constraints, leading to suboptimal conditional generations.
@@ G.6 Different PLM : BERT & T5
Model PPL MAUVE VS(ngram) vs(emb) self-bleu distinct-1 distinct-2
RoBERTa 109.3 0.811 4.417 3.311 0.256 0.407 0.810
BERT 69.5 0.773 4.755 3.659 0.126 0.475 0.834
T5 408.1 0.378 2.256 1.666 0.750 0.415 0.773
Table 15: Performance comparison of T5, BERT, and RoBERTa.
@@ Appendix I Well-Generated Output Examples
Appendix K Details on CDLMs
K.1 Experimental Details
For the case of Diffuseq and Dinoiser Ye et al. (2024), we followed the official repositories to reproduce the results. Results were sampled multiple times with different seeds to evaluate the diversity. Some deviations are as follows. For max-length, we choose 64 for Paradetox, QG, and QQP, and 48 for Deontology. The values were chosen after examining the training set. As for batch size, we followed the original repositories if the parameter was provided. If not, the batch size was chosen using linear interpolation with the size of the training set. Note that unlike other benchmarks, we experimented with Diffuseq-v2 (Gong et al., 2023b) in translation task for a broader comparison with existing baselines.
K.2 Results Interpretations
Examining the results of Diffuseq, it is evident that the grammar score is comparatively lower than that of other models. This outcome is expected, as the outputs from Diffuseq frequently display inaccurate sentence structures, including duplications of words or phrases. Conversely, the outputs from Dinoiser achieve moderate grammar scores but show limited diversity. This finding, coupled with our additional experiments concerning the beam size during Dinoiser generation, suggests that Dinoiser’s performance predominantly relies on memorization.
Appendix L Details on Translation Results
L.1 Datasets & Observations
@@
L.2 Comparison Between Easy-First and Our Proposed Method
@@
L.3 Experimental Details
NAR Transformer & CMLM We utilized the official repository to produce obtain the results, with the default architecture, optimization, and inference configurations. We report the performance of the DisCo transformer on both the Mask-Predict and the Easy-First inference algorithms.
Diffuseq-v2 For Diffuseq-v2, we employ the vocabs of mBERT and choose 128 as max length for ende translation. Other settings are identical as in Appendix K.1.
SEDD The SEDDLou et al. (2024) model, originally designed for open-ended text generation, was adapted in this study to facilitate conditional generation. To align the model’s architecture with the specific requirements of the structured dataset, several modifications were implemented in both hyperparameters and preprocessing protocols. Specifically, the input and output token lengths were constrained to a range of 64 to 128 tokens, ensuring a more appropriate fit to the dataset’s structural characteristics. Moreover, distinct special tokens were introduced to clearly differentiate between input and output sequences, thereby enhancing the model’s ability to distinguish between these components during training. Individual data entries were further demarcated by an EOS token to delineate discrete sequences within the training process.
@@
L.4 Experimental Results
L.4.1 NAR Transformer, DisCo
@@
To account for the relatively small train set to valid/test set ratio of the dataset used in our translation experiments, which resulted in a high percentage of <UNK> tokens in the valid/test sets, we also provide results using the dictionary of a pre-trained mBART model Liu (2020). The performance benefits slightly from this change, but still lags behind those of other models.
L.4.2 Diffuseq-v2
@@
For Diffuseq-v2, we conducted additional experiments using the same model trained on Paradetox. We observed that the entropy of token prediction probabilities in the translation model was orders of magnitude higher, indicating a greater level of uncertainty in its predictions. Similarly, the ratio of the nearest token distance to the average distance of the top five nearest tokens was significantly larger in the translation model. This analysis suggests that a simple rounding approach from continuous to discrete space may be insufficient for machine translation, at least in low-resource settings.

### v5

@@ 3.2 DDLM with Entropy-based Denoising
where $p_{\theta}(x^{\prime}_{i};X^{(t)})$ is the softmax probability of token $x^{\prime}_{i}$ at position $i$ in sequence $X^{(t)}$, and $t$ denotes the diffusion timestep. We then select high-entropy positions for updating:
@@ 5.1 Tasks & Details
We conduct experiments on various conditional generation datasets. Detailed explanation of the conditional generation datasets are in Appendix F.1. In particular, we focus on two datasets that significantly differ in their level of conditional constraints: RocStories (Mostafazadeh et al., 2016), which is relatively open-ended, and Paradetox (Logacheva et al., 2022), which imposes the strongest conditional constraints. We select the conditional dataset that GPT-2 faces in generating sentences of appropriate length under specified conditional constraints. The maximum lengths of Paradetox and RocStories is set to 64, based on data statistics, and other details are in Appendix F. We test 20 conditions with 5 outputs in total 100, which is not used for training. The number of steps of our model is configured to 5 with a naive categorical sampling with a sample size of 20 and select final 5 samples based on Perplexity score. We use 1 A100 GPU with the batch size as 256.
@@ 5.2 Baselines
For ARMs (Vaswani et al., 2023), we employ GPT-2 (Radford et al., 2019) and gpt-3.5-turbo***https://platform.openai.com/docs/models/gpt-3-5 with four-shot prompt. More experimental details of GPT-3.5 can be found in Appendix J. For CMLMs, we utilize CMLM-Mask-Predict (Ghazvininejad et al., 2019a) and DisCo-Easy-First (Kasai et al., 2020), which are transformer-based NAR models. For CDLMs, our baseline includes DiffuSeq (Gong et al., 2023a), LD4LG (Lovelace et al., 2023), and DINOISER (Ye et al., 2024). DiffuSeq and DINOISER is designed for sequence-to-sequence applications, and LD4LG adopt BART as denoising init point. For DDLMs, we utilize DiffusionBERT (He et al., 2022), applying BERT into DDLMs, and SEDD (Lou et al., 2024), a powerful open-ended generation DDLM. For SEDD, we download the pre-trained version and fine-tune it. More details are in Appendix F.3.
@@ 5.3 Metrics
In addition to our theoretically guided methods, we evaluate performance using multiple metrics. Specifically, we use Perplexity (PPL) based-on GPT-2 Large and GPT-2 XL as an automated metric, MAUVE Pillutla et al. (2021) to assess style consistency between the training data and generated outputs, SOME (Yoshimura et al., 2020) to score the grammar, Mean Opinion Score (MOS) from human evaluations to gauge text quality, and LLM score such as LLM-c (Lin and Chen, 2023) to measure the plausibility of the narratives as a sub-metric.
@@
Following our quality assessment, we evaluate diversity through three different measures: an automatic frequency-based metric n-gram Vendi Score(VS n-gram) Friedman and Dieng (2023), a neural network–based semantic metric SimCSE Vendi Score (VS emb), and a human evaluation score MOS. The detailed descriptions of metrics are provided in Appendix F.3 and H.1.
@@ 6 Results
Overall, Diffusion-EAGS consistently demonstrates strong performance across diverse conditional generation tasks, combining low perplexity and high human evaluation scores with the ability to generate richly varied text. Detailed results are in Appendix G and examples are in Appendix I.
@@ F.2 Dataset Explanations
Deontology The objective of Deontology (Hendrycks et al., 2023) is to evaluate the capability of models to make ethical judgments from a deontological perspective. The dataset contains scenarios focusing on interpersonal dynamics and everyday occurrences.
@@ F.3 Experimental Details
We employ roberta-base as MLM with learning rate 5e-4. The maximum lengths for QG, QQP, and Paradetox is set to 64, while for Deontology and DialogSum set to 48 and 292, respectively, based on data statistics. We test 20 conditions with 5 outputs in total 100, which is not used for training. The number of steps is configured to 5. We then perform a naive categorical sampling with a sample size of 20 and select final 5 samples based on PPL. We use 1 A100 GPU with the batch size as 256.
For the case of ARMs, CMLMs, CDLMs, and DDLMs, we follow the official repositories to reproduce the results. Results are sampled multiple times with different seeds to evaluate the diversity. For hyperparameters, we follow the original repositories if the parameter is provided, except for modifying the number of samples as 5 and max_length parameters according to data statistics. Note that unlike other benchmarks, we experiment with Diffuseq-v2 (Gong et al., 2023b) in translation task for a broader comparison with existing baselines. Moreover, experimental details of LLMs are in Appendix J, machine translation in Appendix H.
@@
To measure the quality of the generated texts, we use Perplexity based-on GPT-2 Large and GPT-2 XL, SOME (Yoshimura et al., 2020), the grammar metric based on corpus, LLM-c (Lin and Chen, 2023) to measure the plausibility of the narratives, LLM-t (Koh et al., 2024a) to measure toxicity, and MAUVE Pillutla et al. (2021), measuring a reflectiveness of training dataset characteristics of generate outputs. MAUVE score of 1 indicates that the output perfectly matches the training dataset as a neural database. For Mean Opinion Score (MOS), we get 5 outputs from each condition. For a fair MOS comparison, if GPT-3.5-turbo refuses to provide an answer or if sentence completeness is compromised by condition consisting of “rtttt,” or extreme elliptical expressions, we exclude such relevant condition from our evaluation target. Subsequently, hired four integrated ph.d course work annotators in the university NLP research lab evaluate the generated text based on two criteria: (1) semantic reflectiveness of the condition, indicating how accurately the condition is represented in the text, and (2) sentence completeness, assessing overall grammatical and semantic coherence. Each criterion was rated on a scale from 0 to 1. Subsequently, these scores are normalized and averaged to obtain a final score ranging from 0 to 1. In our evaluation, Fleiss’ kappa Fleiss (1971) is exceeded 0.7 as assessing sentence quality is both intuitive and relatively non-controversial among the annotators.
@@
Ours 0.409 0.174 0.536 4.114 2.591 0.252 0.253 0.632
SEDD 0.179 0.032 0.999 4.216 2.576 0.211 0.200 0.609
DINOISER 0.209 0.031 0.337 1.247 1.227 0.926 0.256 0.633
Table 10: DialogueSum Experiment
@@ G.1 Fine-Grained Comparison
Diffusion-EAGS demonstrates the highest MAUVE score in Table 8-ParaDetox, and high level of text quality surpassing that of GPT-2 in Table 9 in text quality. ParaDetox is colloquial dataset including slang, numerous abbreviations, and various perturbations, so our model demonstrate robustness to such perturbations. As for diversity, our model consistently outperforms GPT models in VS(ngram) and VS(emb) in Table 28, and 9.
Notably, CDLMs demonstrate a noticeable deficiency in diversity. Examining the results of Diffuseq, it is evident that the grammar score is comparatively lower than that of other models. This outcome is expected, as the outputs from Diffuseq frequently display inaccurate sentence structures, including duplications of words or phrases. Conversely, the outputs from Dinoiser achieve moderate grammar scores but show limited diversity. This finding, coupled with our additional experiments concerning the beam size during Dinoiser generation, suggests that Dinoiser’s performance predominantly relies on memorization. In contrast, our model excels at producing significantly more diverse sequences. Furthermore, our models require only a few steps, while resulting in higher quality and diversity.
@@ G.2 Quality Recheck – LLM score & Dialogue Summarization
For Question Generation (QG), we employ the widely adopted GPT-as-a-Judge framework Zheng et al. (2023) to evaluate the quality of generations produced by our model and the baselines on the QG dataset. We adopt a pairwise evaluation setting, following the system and input prompts specified in Zheng et al. (2023) for the pairwise comparison. The factors specified to be evaluated are 1) coherency 2) grammatical correctness 3) semantic soundness 4) diversity and 5) being a more reasonable question to the input (condition) text. We employ the gpt-4 model. The result is in Table 13.
@@
Below, we report the Mean Opinion Score (MOS) averages and standard deviations (std) in the following order: DiffusionBERT, LD4LG, GPT-2, Dinoiser, and our method. First, the average scores of semantic reflection are 0.98, 0.90, 0.94, 0.98, and 0.97, respectively, with standard deviations of 0.14, 0.30, 0.24, 0.14, and 0.16. Second, the average scores of sentence completeness are 0.78, 0.92, 0.72, 0.84, and 0.90, respectively, with standard deviations of 0.18, 0.14, 0.28, 0.15, and 0.15. Third, average scores of diversity are 2, 1, 2.65, 1, and 4.6, respectively, with standard deviations of 1.3, 0, 1.45, 0, and 0.7. GPT-3.5-turbo’s std is 0 for quality MOS and 0.83 for diversity MOS.
@@ G.3 Machine Translation : Bilinguality & Low Resource Settings
Labeled datasets used in conditional generation tasks are typically limited in size and sometimes multilingual. To further assess our model’s performance in conditional generation, particularly in terms of language extension and resource scarcity, we conduct additional experiments on a translation task. We conduct additional experiments on CMLMs such as Mask-and-Predict and Easy-First, diffusion models such as Diffuseq-v2 (Gong et al., 2023b) and SEDD, traditional translation models such as mBART-50 (Tang et al., 2020) and NLLB. For evaluation metrics, we utilize sacreBLEU (Post, 2018) and neural-net scores such as COMET (Rei et al., 2020) and XCOMET (Guerreiro et al., 2023). More details are provided in Appendix K.
@@ G.4 Diversity Analysis
We summarize the generation trends of the models presented in Table below. We observe that when a fine-tuned GPT-2 is tasked with strongly constrained conditional generation, it struggles to properly terminate sentences with an <eos> token. In particular, it shows limitations when handling semantic leaps or clearly delineated structural constraints, leading to suboptimal conditional generations.
@@
Model PPL MAUVE VS(ngram) vs(emb) self-bleu distinct-1 distinct-2
RoBERTa 109.3 0.811 4.417 3.311 0.256 0.407 0.810
BERT 69.5 0.773 4.755 3.659 0.126 0.475 0.834
T5 408.1 0.378 2.256 1.666 0.750 0.415 0.773
Table 15: Performance comparison of T5, BERT, and RoBERTa.
@@ Appendix I Well-Generated Output Examples
Appendix K Details on Translation Results
K.1 Datasets & Observations
@@
K.2 Comparison Between Easy-First and Our Proposed Method
@@
K.3 Experimental Details
NAR Transformer & CMLM We utilize the official repository to produce obtain the results, with the default architecture, optimization, and inference configurations. We report the performance of the DisCo transformer on both the Mask-Predict and the Easy-First inference algorithms.
Diffuseq-v2 For Diffuseq-v2, we employ the vocabs of mBERT and choose 128 as max length for ende translation. Other settings are identical as official repository.
SEDD The SEDDLou et al. (2024) model, originally designed for open-ended text generation, is adapted in this study to facilitate conditional generation. To align the model’s architecture with the specific requirements of the structured dataset, several modifications are implemented in both hyperparameters and preprocessing protocols. Specifically, the input and output token lengths are constrained to a range of 64 to 128 tokens, ensuring a more appropriate fit to the dataset’s structural characteristics. Moreover, distinct special tokens are introduced to clearly differentiate between input and output sequences, thereby enhancing the model’s ability to distinguish between these components during training. Individual data entries are further demarcated by an EOS token to delineate discrete sequences within the training process.
@@
K.4 Experimental Results
K.4.1 NAR Transformer, DisCo
@@
To account for the relatively small train set to valid/test set ratio of the dataset used in our translation experiments, which results in a high percentage of <UNK> tokens in the valid/test sets, we also provide results using the dictionary of a pre-trained mBART model Liu (2020). The performance benefits slightly from this change, but still lags behind those of other models.
K.4.2 Diffuseq-v2
@@
For Diffuseq-v2, we conduct additional experiments using the same model trained on Paradetox. We observe that the entropy of token prediction probabilities in the translation model is orders of magnitude higher, indicating a greater level of uncertainty in its predictions. Similarly, the ratio of the nearest token distance to the average distance of the top five nearest tokens is significantly larger in the translation model. This analysis suggests that a simple rounding approach from continuous to discrete space may be insufficient for machine translation, at least in low-resource settings.